Objectives To make basic measurements of length, mass and time To practice graphing using spreadsheet software and paper and pencil. 1. Lenght Estimate the length of a meter by putting a pen or pencil at one end of a table and placing a second pen or pencil about one meter away from the first. Use the meter tape measure to measure the actual length of your meter estimate. Record the length of your meter estimate Calculate the percent error of your estimated meter from an actual meter. 2. Time Estimate a 30-s time period while someone else times you using a stopwatch. (If you do not have a partner, you can do this experiment by closing your eyes; start the stopwatch and stop it when you think 30 s have elapsed.) Record the actual time of your estimate. Calculate the percent error of the estimate from the actual time. 3. Mass Pick up a small paperback book or similar small object and estimate its mass. Determine the actual mass of the object using the 500-g spring scale. Record the estimated mass and the actual mass and calculate the percent error. Data Table 1: Estimation of various measurements Measurement Length (m) Time (s) Mass (kg) Estimated Actual | % Error
Answer #1

1.

Le = estimated length = 0.950 m

L = actual length = 1 m

percentage error is given as

%age error = (Le - L) x 100 /L

%age error = (0.950 - 1) x 100 /1

%age error = - 5 %

so error is by 5%

2.

Te = estimated time = 35 sec

T = actual time = 30 sec

percentage error is given as

%age error = (Te - T) x 100 /T

%age error = (35 - 30) x 100 /30

%age error =16.67 %

so error is by 16.67%

3.

me = estimated mass = 450 g

m = actual mass = 500 g

percentage error is given as

%age error = (me - m) x 100 /m

%age error = (450 - 500) x 100 /500

%age error =- 10 %

so error is by 10%